# Diversifying halo structures in two-component self-interacting dark matter models via mass segregation

### v1

@@ Abstract
Introduction. While the standard cold dark matter (CDM) paradigm has been well-tested at large scales, it faces challenges at small scales, particularly in accounting for the diversity of observed galactic rotation curves and inner halo densities Bullock and Boylan-Kolchin (2017); Salucci (2019); Gentile et al. (2004); van Dokkum et al. (2022); Bonaca et al. (2018). One promising solution is self-interacting dark matter (SIDM), where dark matter particles interact through elastic scatterings, leading to gravothermal evolution that results in both cored or cuspy density profiles in halos Tulin and Yu (2018); Adhikari et al. (2022); Spergel and Steinhardt (2000); Kochanek and White (2000); Kamada et al. (2017); Ren et al. (2019); Santos-Santos et al. (2020); Correa et al. (2022); Yang et al. (2023); Zavala et al. (2019); Kaplinghat et al. (2019); Turner et al. (2021); Slone et al. (2023); Silverman et al. (2022); Yang and Yu (2022); Yang et al. (2024a); Yang (2024); Hou et al. (2025). Recent studies Nadler et al. (2023); Kong et al. (2022); Zhang et al. (2024, 2025a, 2025b); Sameie et al. (2020); Yang et al. (2020); Kong and Yu (2025); Balberg and Shapiro (2002); Pollack et al. (2015); Choquette et al. (2019); Feng et al. (2022); Meshveliani et al. (2023); Alonso-Álvarez et al. (2024); Jiang et al. (2025) further show that it may explain a variety of extreme observations, including dark matter deficient galaxies, supermassive black holes at high redshifts, strong lensing perturbers, excessive small-scale lenses, and even the final parsec problem of binary black hole mergers. However, in minimal SIDM scenarios involving a single dark matter species coupled to a light mediator, efficient annihilation channels typically lead to signals in conflict with Cosmic Microwave Background (CMB) and other indirect detection constraints Bringmann et al. (2017). This tension motivates the exploration of more complex scenarios, such as asymmetric dark matter models Kaplan et al. (2009); Petraki et al. (2012). These models propose a dark sector with two particle species exhibiting an asymmetry, analogous to the matter-antimatter asymmetry in the visible sector Kaplan et al. (2009); Petraki et al. (2012); Cyr-Racine and Sigurdson (2013); Petraki and Volkas (2013); Petraki et al. (2014); Zurek (2014); Han et al. (2024). In the early universe, the symmetric components efficiently annihilate, leaving behind the asymmetric population that sets the relic dark matter abundance. Since the relic density is no longer tied to the annihilation cross section, these models allow greater flexibility to realize strong self-interactions. Moreover, just as the proton and electron have different masses but interact through the Coulomb force, the two dark matter components can have different masses and interact among themselves Petraki et al. (2014); Cyr-Racine and Sigurdson (2013).
The introduction of a light mediator naturally gives rise to a velocity-dependent SIDM cross section, a framework that has been extensively studied in recent years and has shown promise in addressing a range of anomalous observations Valli and Yu (2018); Zavala et al. (2019); Correa (2021); Ebisu et al. (2022); Silverman et al. (2022); Feng et al. (2009, 2010); Buckley and Fox (2010); Loeb and Weiner (2011); Tulin et al. (2013); Cyr-Racine et al. (2016); Kaplinghat et al. (2016); Agrawal et al. (2017); Vogelsberger et al. (2012, 2016); Robertson et al. (2019); Nadler et al. (2020); Turner et al. (2021); Bhattacharyya et al. (2022); Zeng et al. (2022); Robertson et al. (2017); Kahlhoefer et al. (2014); Fischer et al. (2021, 2022); Nadler et al. (2025). These studies typically assume either a single dark matter component that is its own antiparticle (i.e., a Majorana particle) or a two-component scenario where both components have equal or nearly equal masses. In this work, we present cosmological $N$-body simulations of a two-component SIDM model with a $3:1$ mass ratio. We show that interactions between the two species lead to pronounced mass segregation, modifying the range of inner density profiles compared to single-component SIDM. Previous simulations involving two-component dark matter have focused on either large mass ratios, as in atomic dark matter, or near-degenerate states Low et al. (2025); Roy et al. (2023); Schutz and Slatyer (2015), but neither has considered the effects of gravothermal evolution. Our framework, though simple, is broadly applicable and introduces a more versatile mechanism that generates novel structural features. These may help resolve existing tensions and ongoing debates stemming from recent precision constraints on SIDM from isolated galaxies, bullet clusters, halo shapes, and the Tully-Fisher relation Correa et al. (2025); Kong et al. (2024); Yang et al. (2024b); Ebisu et al. (2022); Vargya et al. (2022); Correa (2021).
@@
The interplay between mass segregation and self-interactions thus broadens the phenomenology of SIDM. In particular, segregation allows some halos to exhibit core-collapsing features even at modest cross sections where one-component models would form overly large cores. The extent of this effect depends on multiple factors: the mass difference and cross-component scattering rates determine the degree of mass segregation, while the self-interaction cross section of the heavier component regulates the diversity in density profiles. These effects intertwine, collectively reshaping small-scale structures in the inner halo regions. In the limit where the two particle species have equal masses, mass segregation disappears, and the two-component model effectively reduces to the one-component SIDM scenario, although cross-species interactions remain governed by a Rutherford-like scattering cross section.
@@ Acknowledgements.
The authors were supported in part by the National Key Research and Development Program of China (No. 2022YFF0503304), the Project for Young Scientists in Basic Research of the Chinese Academy of Sciences (No. YSBR-092). Software: Rockstar Behroozi et al. (2013), Scipy Virtanen et al. (2020), NumPy van der Walt et al. (2011), Matplotlib Hunter (2007).
@@ Appendix A Implementation of differential scatterings
$\chi_{H}-\chi_{H}$ $m_{H}=\frac{3}{2}m_{0}$ $\frac{\sigma_{H}}{m_{H}}=\frac{3}{2}\frac{\sigma_{M}}{m_{0}}$ $w_{1}=\frac{3}{2}w_{M}$
@@ Appendix B Fitting the density profiles
We find that the parametric profile commonly used to model one-component SIDM halos Yang et al. (2024a); Yang (2024); Yang et al. (2025); Hou et al. (2025) can still be used to describe halos in two-component dark matter. Specifically, we fit the density profiles using the following $\beta 4$ profile:

### v2

@@ Abstract
Introduction. While the cold dark matter (CDM) paradigm has been remarkably successful in explaining the large-scale structure of the universe through gravitational interactions, the microscopic nature of dark matter, from its composition to interaction properties, remains fundamentally unknown. Uncovering its particle nature requires physics beyond the Standard Model and provides a unique window into early-universe processes, including the origin of the universe, baryogenesis, and the interplay between inflation and the emergence of cosmic structure Cirelli et al. (2024); Arbey and Mahmoudi (2021); Balazs et al. (2024). Among the various theoretical frameworks, self-interacting dark matter (SIDM) has emerged as a compelling candidate, particularly in addressing persistent discrepancies on galactic and sub-galactic scales Bullock and Boylan-Kolchin (2017); Salucci (2019); Gentile et al. (2004); van Dokkum et al. (2022); Bonaca et al. (2018). Elastic scattering among dark matter particles in SIDM models drives gravothermal evolution, leading to either cored or cuspy halo density profiles, depending on the interaction strength and the system’s evolutionary stage Tulin and Yu (2018); Adhikari et al. (2022); Spergel and Steinhardt (2000); Kochanek and White (2000); Kamada et al. (2017); Ren et al. (2019); Santos-Santos et al. (2020); Correa et al. (2022); Yang et al. (2023); Zavala et al. (2019); Kaplinghat et al. (2019); Turner et al. (2021); Slone et al. (2023); Silverman et al. (2022); Yang and Yu (2022); Yang et al. (2024a); Yang (2024); Hou et al. (2025); Robertson et al. (2019); Chu et al. (2019, 2020); Fischer et al. (2021, 2022); Zeng et al. (2022); Nadler et al. (2025). This mechanism provides a natural explanation for a variety of extreme and otherwise puzzling astrophysical observations, including dark matter-deficient galaxies, supermassive black holes at high redshifts, strong lensing perturbers, excessive small-scale lenses, and even the final-parsec problem in binary black hole mergers Nadler et al. (2023); Kong et al. (2022); Zhang et al. (2024, 2025a, 2025b); Sameie et al. (2020); Yang et al. (2020); Kong and Yu (2025); Balberg and Shapiro (2002); Pollack et al. (2015); Choquette et al. (2019); Feng et al. (2022); Meshveliani et al. (2023); Alonso-Álvarez et al. (2024); Jiang et al. (2025).
Although the non-luminous nature of dark matter renders its internal composition unobservable in principle, this degeneracy can be lifted in multi-component scenarios with cross-component interactions, as we demonstrate in this work. The possibility of multiple dark matter species is not only a natural extension but is also well-motivated theoretically. In minimal SIDM models involving a single species coupled to a light mediator, efficient annihilation channels often produce signals in conflict with Cosmic Microwave Background and indirect detection constraints Bringmann et al. (2017). These tensions motivate consideration of richer dark sector structures, such as asymmetric dark matter models Kaplan et al. (2009); Petraki et al. (2012). These models postulate two dark matter species with an asymmetry analogous to the baryon asymmetry in the visible sector Kaplan et al. (2009); Petraki et al. (2012); Cyr-Racine and Sigurdson (2013); Petraki and Volkas (2013); Petraki et al. (2014); Zurek (2014); Han et al. (2024). In the early universe, the symmetric components annihilate efficiently, leaving behind the asymmetric population that determines the relic abundance. Because the relic density is no longer determined by the annihilation cross section, these models naturally accommodate strong self-interactions. Furthermore, just as the proton and electron differ in mass while interacting via the electromagnetic force, dark matter components can have distinct masses while coupling through a dark-sector interaction Petraki et al. (2014); Cyr-Racine and Sigurdson (2013). This setup introduces novel structure formation dynamics, which we explore via cosmological $N$-body simulations of two-component SIDM. In particular, we focus on a mass ratio of $3:1$ between the species and demonstrate that inter-species interactions lead to significant mass segregation, resulting in a broader range of inner density profiles compared to single-component SIDM.
Prior simulations of two-component dark matter have typically focused on either large mass ratios, as in atomic dark matter, or near-degenerate states Low et al. (2025); Roy et al. (2023); Schutz and Slatyer (2015), without considering the effects of gravothermal evolution. Ref. Choquette et al. (2019) performed N-body simulations of a dark matter subcomponent that self-interacts, focusing on its gravothermal collapse as a potential seed for a supermassive black hole. In contrast, our work focuses on the density profiles of entire dark matter halos, demonstrating how two-component SIDM reshapes their structural diversity. Despite its simplicity, our framework is broadly applicable and offers a versatile mechanism for generating either diffuse or dense halo structures. These features may help resolve outstanding tensions in SIDM interpretations of recent precision data, including observations of small-scale lenses, bullet clusters, halo ellipticities, and the Tully-Fisher relation Correa et al. (2025); Kong et al. (2024); Yang et al. (2024b); Ebisu et al. (2022); Vargya et al. (2022); Correa (2021).
@@
The interplay between mass segregation and self-interactions thus broadens the phenomenology of SIDM. In particular, segregation allows some halos to exhibit core-collapsing features even at modest cross sections where one-component models would form overly large cores. This behavior can be traced to the maintenance of pressure equilibrium during cross-component scatterings. A decrease in the velocity of the heavier component, $v_{H}$, must be compensated by an increase in its density to keep the pressure, $\propto\rho_{H}v_{H}^{2}$, constant.The extent of this effect depends on multiple factors: the mass difference and cross-component scattering rates determine the degree of mass segregation, while the self-interaction cross section of the heavier component regulates the diversity in density profiles. These effects intertwine, collectively reshaping small-scale structures in the inner halo regions. In the limit where the two particle species have equal masses, mass segregation disappears, and the two-component model effectively reduces to the one-component SIDM scenario, although cross-species interactions remain governed by a Rutherford-like scattering cross section.
@@
Note added: After the submission of this work, Ref. Patil and Fischer (2025) explored halo evolution under cross-component interaction with mass segregation. Ref. Yang et al. (2025a) further elaborated on observational signatures from SIDM with mass segregation in dwarf halos and cluster substructures. The findings of these works align and complement.
@@ Acknowledgements.
We thank Hai-Bo Yu, Haipeng An, and Xiaoyong Chu for helpful discussion. The authors were supported in part by the National Key Research and Development Program of China (No. 2022YFF0503304), the Project for Young Scientists in Basic Research of the Chinese Academy of Sciences (No. YSBR-092). Software: Rockstar Behroozi et al. (2013), Scipy Virtanen et al. (2020), NumPy van der Walt et al. (2011), Matplotlib Hunter (2007).
@@ Appendix A Implementation of differential scatterings
$\chi_{H}-\chi_{H}$ $m_{H}=\frac{3}{2}m_{0}$ $\frac{\sigma_{H}}{m_{H}}=\frac{3}{2}\frac{\sigma_{M}}{m_{0}}$ $w_{1}=\frac{2}{3}w_{M}$
@@ Appendix B Fitting the density profiles
We find that the parametric profile commonly used to model one-component SIDM halos Yang et al. (2024a); Yang (2024); Yang et al. (2025b); Hou et al. (2025) can still be used to describe halos in two-component dark matter. Specifically, we fit the density profiles using the following $\beta 4$ profile: